NEW INFRARED TRANSITIONS OF SOLID PARAHYDROGEN IN THE SPECTRAL REGION 7900 - 9900 cm^{−1} \underline{M. MENGEL}, B. P. WINNEWISSER AND M. WINNEWISSER, Physikalisch-Chemisches Institut, Justus-Liebig-Universität, Heinrich-Buff-Ring 58, D-35392 Giessen. In our most recent measurements of infrared transitions of solid parahydrogen with low ortho impurities, using a White-type external multireflection system ^{a} for a sample path length of 38 cm, we have detected new extremely weak features in the spectral region 7900 - 9900 cm^{−1}. Among them are the single transition

• Y_{1←0}(0) at 7991.85 (1) cm^{−1} and the double transitions

• Q_{1←0}(0) + Q_{1←0}(0) at 8306.0 (1) cm^{−1}

• U_{1←0}(0) + Q_{1←0}(1) at 9407.67 (5) cm^{−1}

• Q_{1←0}(0) + S_{1←0}(0) at 9747.8 (1) cm^{−1}

• U_{1←0}(1) + Q_{1←0}(0) at 9837.34 (5) cm^{−1}. The Y_{1←0}(0) transition is the second transition with ΔJ = 8 observed in a molecular spectrum after the Y_{0←0}(0) transition ^{b} and is even weaker than the latter by an appreciable factor. Intensity parameters will be reported for all of these transitions. The data were taken at a sample temperature of 12.3 K.

aR. Steinhoff, K. V. S. R. M. Winnewisser, Appl. Optics 32, 6577-6581 (1993).

bR. A. Steinhoff, B. P. Winnewisser and M. Winnewisser, Phys. Rev. Lett. 73, 2833 - 2836 (1994).